# Pencils of plane cubics with one base point

### v1

@@ 1. Introduction
The key point of the construction of $|{\mathcal{V}}|$ is the fact that the base point necessarily is a $9$-torsion point which is not $3$-torsion of a smooth member (once the origin has been fixed in a flex point). The theory of points of order nine on a cubic has a long and glorious history, dating back more than 150 years with the works of Cayley [7], Hart [11] and Halphen [10], Salmon [15]. More recently, the same kind of topics were considered by Remo Gattazzo in [9], together with a modern viewpoint on the beautiful construction of the so-called tangential triangles of a cubic 111“It may be proposed to determine the points at which it shall be possible to draw cubics osculating the given curve in nine points; or in other words, such that the third tangential D may coincide with the original point A.” Salmon, [15].. To the present days, the problem was considered for instance in [2]. We give in Section 5 an overview of Gattazzo’s results in relation to ours. Using both these techniques and our approach we can produce explicit examples of equations for $|{\mathcal{V}}|$: see Example 2.4 and 5.4.
@@
In Section 9 we see that all fibrations associated to pencils satisfying (C1) and (C2) are rational extremal elliptic fibrations (see Definition 9.1), and we show where they appear in Miranda-Persson’s classification [13, Theorem 4.1].
@@ Acknowledgements
We would like to thank Alice Garbagnati, for explaining us the classification of Miranda and Persson. We owe her Section 9. We also thank Claudia Alcàntara, Ciro Ciliberto and Letterio Gatto for their interest in our work and their kind encouragement. R.M. is partially supported by the PRIN project 2022L34E7W “Moduli Spaces and Birational Geometry”. G.P.P. and L.S. are partially supported by the PRIN project 20228JRCYB “Moduli spaces and special varieties”. All the authors are members of the GNSAGA - INdAM.
@@ 3. Uniqueness of the construction
We have seen that any curve can be an element of a pencil satisfying the assumptions of 1.1, thus answering to point $(i)$ of 1.2: the codimension is in this case 00. We now see that any pencil of degree $3$ satisfying the conditions of 1.1 can be realised with the construction of Proposition 2.2. Let us first recall the following elementary result on the local intersection of plane curves, which we will repeatedly use in the paper. For a proof see for instance [14, Lemma 1.3.8].
@@ 6. Uniqueness up to projective transformations
Note that by Proposition 3.3, although the equations are quite different, the pencils of Example 5.4 and 2.4 are the same, up to projectivity.
@@ 8. Pencils with one base point which is a flex
With the same arguments as in Lemmas 2.3, 4.1 and 4.2 we can prove the following:
@@ 9. Comparison with Miranda-Persson classification
The fibrations $f$, $g$ and $g^{\prime}$ are all rational extremal elliptic fibrations. They correspond in Miranda-Persson’s list in [13, Theorem 4.1] to the elements $X_{1119}$, $X_{211}$ and $X_{22}$, respectively.
@@
Now we prove that all our fibrations are extremal. Recall the formula of Shioda-Tate (see [13, Corollary 2.4]), holding for any elliptic fibration $h\colon X\to B$:
@@
We can see the isotriviality of the fibration $g^{\prime}$ also in the list of [13, Theorem 4.1], because the modular map $J\colon{\mathbb{P}}^{1}\to\overline{\mathcal{M}}_{1,1}\cong{\mathbb{P}}^{1}$ has degree 00. On the other hand, the modular map for the fibration $f$ corresponding to $|{\mathcal{V}}|$ has degree $12$, as proved in loc. cit. In this case, the three sections that compose the Mordell-Weil group $\Phi(X)$ identify a tangential triangle. Given a smooth plane cubic $C$ and fixing a point $p$ of type $9$ on $C$, there exists a fibre $F_{t}$ of $f$ with the $3$ points $\Phi(X)\cap F_{t}$ corresponding to the tangential triangle associated to $p$. With this choice $[(C,p)]\in\mathcal{M}_{1,1}$ is the image of the modular map corresponding to the fibre $F_{t}$. It seems that the preimage $J^{-1}([(C,p)]$ should identify a subset of 12 tangential triangles on $C$. So, there seems to be a natural partition of the tangential triangles into two subsets of cardinality 12 each, that might be further investigated.

### v2

@@ 1. Introduction
The key point of the construction of $|{\mathcal{V}}|$ is the fact that the base point necessarily is a $9$-torsion point which is not $3$-torsion of a smooth member (once the origin has been fixed in a flex point). The theory of points of order nine on a cubic has a long and glorious history, dating back more than 150 years with the works of Cayley [7], Fano [10], Hart [13], Halphen [12] and Salmon [22]. More recently, the same kind of topics were considered by Remo Gattazzo in [11], together with a modern viewpoint on the beautiful construction of the so-called tangential triangles of a cubic 111“It may be proposed to determine the points at which it shall be possible to draw cubics osculating the given curve in nine points; or in other words, such that the third tangential D may coincide with the original point A.” Salmon, [22]. See also [23, 24 pag.30].. To the present days, the problem is considered for instance in [2] and in [18]. In the recent preprint [15] the same pencil is constructed in Example 46. We give in Section 5 an overview of Gattazzo’s results in relation to ours. Using both these techniques and our approach we can produce explicit examples of equations for $|{\mathcal{V}}|$: see Example 2.4 and 5.4.
@@
In Section 9 we see that all fibrations associated to pencils satisfying (C1) and (C2) are rational extremal elliptic fibrations (see Definition 9.1), and we show where they appear in Miranda-Persson’s classification [19, Theorem 4.1].
@@ Acknowledgements
We are grateful to Alice Garbagnati, for explaining us the classification of Miranda and Persson. We owe her Section 9. We also thank Claudia Alcàntara, Ciro Ciliberto and Letterio Gatto for their interest in our work and their kind encouragement. We are pleased to thank János Kollár for suggesting Remark 6.2 related to his recent preprint [15], and giving us some interesting references Remark 7.2.
R.M. is partially supported by the PRIN project 2022L34E7W “Moduli Spaces and Birational Geometry”. G.P.P. and L.S. are partially supported by the PRIN project 20228JRCYB “Moduli spaces and special varieties”. All the authors are members of the GNSAGA - INdAM.
@@ 3. Uniqueness of the construction
We have seen that any curve can be an element of a pencil satisfying the assumptions of 1.1, thus answering to point $(i)$ of 1.2: the codimension is in this case 00. We now see that any pencil of degree $3$ satisfying the conditions of 1.1 can be realised with the construction of Proposition 2.2. Let us first recall the following elementary result on the local intersection of plane curves, which we will repeatedly use in the paper. For a proof see for instance [20, Lemma 1.3.8].
@@ 4. Singular elements and non-isotriviality
Remark 4.2.
The second part of the proof, that excludes the existence of a cuspidal cubic with the cusp different from the base point $p$, can be proved also with Lemma 42 of [15].
@@ 6. Uniqueness up to projective transformations
An alternative proof of the uniqueness up to projective transformations follows from [15, Example 46]: starting from a nodal cubic $C_{0}$ at $p$, there exists a unique pencil $|V|$ satisfying $(C1)$, $(C2)$ and $(C3)$ with $p$ as base point with all the members tangent to $C_{0}$ on a fixed branch at $p$.
As suggested by János Kollár in a private communication, this can be used to give a classification of this kind of pencils up to projective transformation over any field $k$, possibly of characteristic different from $2,3$. Indeed, up to projective transformations, a nodal cubic on ${\mathbb{P}}^{2}_{k}$ with tangent lines at the node defined over $k$ has equation
$x^{3}+\lambda y^{3}-xyz=0,$ (6.2)
where $\lambda$ is in $k^{\ast}$. Note that the line passing through all the nonsingular flexes over an extension of $k$ is defined on the field $k$ and we obtain Equation (6.2) by imposing that such line has equation $z=0$. Two of these equations associated to $\lambda$ and $\lambda^{\prime}$ gives the same cubic if and only if $\lambda/\lambda^{\prime}$ belongs to $(k^{\ast})^{3}$. As a consequence the pencils are classified by $k^{\ast}/(k^{\ast})^{3}$.
@@ 7. Comparison with Beauville’s pencils
Remark 7.2.
In the context of Enriques surfaces, the correspondence between (6.1) and (F6) is described in [9, Remark 8.9.31], see also [16, (3.7)]. Similar pictures as Figure 2 have been obtained in [26], [21, Sec.6] and in [17, Sec.2.1].
@@ 8. Pencils with one base point which is a flex
With the same arguments as in Lemmas 2.3, 4.1 and 4.3 we can prove the following:
@@ 9. Comparison with Miranda-Persson classification
The fibrations $f$, $g$ and $g^{\prime}$ are all rational extremal elliptic fibrations. They correspond in Miranda-Persson’s list in [19, Theorem 4.1] to the elements $X_{1119}$, $X_{211}$ and $X_{22}$, respectively.
@@
Now we prove that all our fibrations are extremal. Recall the formula of Shioda-Tate (see [19, Corollary 2.4]), holding for any elliptic fibration $h\colon X\to B$:
@@
We can see the isotriviality of the fibration $g^{\prime}$ also in the list of [19, Theorem 4.1], because the modular map $J\colon{\mathbb{P}}^{1}\to\overline{\mathcal{M}}_{1,1}\cong{\mathbb{P}}^{1}$ has degree 00. On the other hand, the modular map for the fibration $f$ corresponding to $|{\mathcal{V}}|$ has degree $12$, as proved in loc. cit. In this case, the three sections that compose the Mordell-Weil group $\Phi(X)$ identify a tangential triangle. Given a smooth plane cubic $C$ and fixing a point $p$ of type $9$ on $C$, there exists a fibre $F_{t}$ of $f$ with the $3$ points $\Phi(X)\cap F_{t}$ corresponding to the tangential triangle associated to $p$. With this choice $[(C,p)]\in\mathcal{M}_{1,1}$ is the image of the modular map corresponding to the fibre $F_{t}$. It seems that the preimage $J^{-1}([(C,p)]$ should identify a subset of 12 tangential triangles on $C$. So, there seems to be a natural partition of the tangential triangles into two subsets of cardinality 12 each, that might be further investigated.